# Guiding Vision-Language Model Selection for Visual Question-Answering Across Tasks, Domains, and Knowledge Types

### v1

@@ Abstract
Visual Question-Answering (VQA) has become a key use-case in several applications to aid user experience, particularly after Vision-Language Models (VLMs) achieving good results in zero-shot inference. But evaluating different VLMs for an application requirement using a standardized framework in practical settings is still challenging. This paper introduces a comprehensive framework for evaluating VLMs tailored to VQA tasks in practical settings. We present a novel dataset derived from established VQA benchmarks, annotated with task types, application domains, and knowledge types, three key practical aspects on which tasks can vary. We also introduce GoEval, a multimodal evaluation metric developed using GPT-4o, achieving a correlation factor of 56.71% with human judgments. Our experiments with ten state-of-the-art VLMs reveals that no single model excelling universally, making appropriate selection a key design decision. Proprietary models such as Gemini-1.5-Pro and GPT-4o-mini generally outperform others, though open-source models like InternVL-2-8B and CogVLM-2-Llama-3-19B demonstrate competitive strengths in specific contexts, while providing additional advantages. This study guides the selection of VLMs based on specific task requirements and resource constraints, and can also be extended to other vision-language tasks111Code and dataset can be found at the following link: https://github.com/neelabhsinha/vlm-selection-tasks-domains-knowledge-type..
@@ 1 Introduction
Visual Question-Answering (VQA) [1] is a task of answering a question about an image correctly. With the evolution of research, this field has been presented with significant challenges in terms of the nature of the problem. For example, question can be about the image directly [13, 39, 14], or outside the scope of the image where external knowledge may be required [25, 31]. The images can be a photograph, a mathematical chart [26, 19], a document’s screenshot [27], or more.
@@
Dedicated methods [39, 14, 16] have long existed to solve different challenges in VQA. But, with the advancement of Vision-Language models (VLMs) [10, 2, 33, 30, 23] in multimodal research, several applications have started adapting them for VQA, and most of them [18, 33, 34, 10] perform better than dedicated methods. This is because by leveraging pre-training [21, 37] on vast multimodal datasets [8, 35, 7, 26, 27, 25], VLMs can effectively generalize across various types of images, including photographs, charts, and documents, making them highly adaptable to multiple VQA tasks without additional modification. Additionally, VLMs can also incorporate external knowledge beyond the image content, enabling more robust responses to questions with broader context in open-domain settings.
@@
With this shift comes the next question: Which VLM to utilize for a given VQA-based requirement? For example, can the model that is the best in knowledge-based VQA also understand mathematical charts better than the others? The complexity of this question stems from two different directions - the number of VLMs that are being proposed these days, and the diverse nature of tasks in VQA that the VLMs are expected to perform. New VLMs are being released almost every month now [23, 22, 2, 34, 10, 9, 36, 15], and they all are different in their model architecture, training data, prompts, training strategy, size, etc. Due to this, they can possess different strengths and weaknesses. Additionally, users often face technical and business constraints like compute or memory limitations, data privacy risk, associated cost of inference, compliance and regulatory risks, or associated costs of inference, which may lead them to favor specific VLMs over others. For example, one may favor Gemini Flash over Pro due to affordability concerns over cost per input and output token. Second, tasks may vary drastically, and in multiple ways. They can range from types such as Chart Question-Answering [26] or Document Understanding [27], to application domains like Science, Politics, or Sports, and the type of knowledge required, encompassing areas like Geographical Information, Mathematical Reasoning, and beyond. For an application that can fall into one or multiple such categories, how to identify best suited VLM? How to compare them meaningfully using a standardized approach? These are gaps in existing literature. Technical reports of VLMs provide benchmarks and comparison, but they are very theoretical and limited.
To bridge this gap, we propose this end-to-end framework which provides a standardized paradigm to evaluate VLMs for visual question-answering in practical settings. First, we develop a evaluation framework on coming up with a structure that allows this comparison over factors that define practical usage. Our framework allows comparison of VLMs under three separate aspects: task type, application domain, and knowledge type. Multiple datasets like VQAv2 [13], OK-VQA [25], ChartQA [26], etc., provide task instances to train and evaluate models based on different task types, but none of them label their instances on other practical aspects. Our framework proposes a dataset derived from these existing benchmarks where tasks are also classified and labeled with their application domains, and type of knowledge required to carry out the task successfully, as shown in Figure 1. Each task can have multiple tags associated with the above three aspects. Further, recent works [17, 24] have also identified that comparing LLM-generated text with traditional evaluation metrics for open-domain QA [17] as well as Natural Language Generation [24] is challenging and not representative with human evaluations. Because these generative models are trained on large datasets, they may produce correct answers that are unfairly evaluated if they differ from the gold-standard labels. This challenge would also transfer to VLMs. To mitigate this, we also propose GoEval, which is a multimodal evaluation metric based on GPT-4o [30], and shows superior correlation with human judgement than previously used metrics when used with VLM generated outputs. Our framework is for VQA, but can be adapted to other vision-language tasks.
If application developers utilize VLMs for VQA, some might prefer openly available VLMs [34, 10, 22] due to compliance and data privacy requirements, some may need small VLMs [34, 10] which can be hosted on-device or using less resources, and some can have other business or economic constraints. Further, how an application plans to use VLMs can also vary. They can be adapted to align closer to downstream requirements [6], in a customized pipeline [12], or directly without any modifications [18]. To accommodate all these needs, we evaluate 10 variations of 8 different VLMs using our method, and identify the best models for different tasks amongst the currently existing VLMs, allowing users to select the best alternative as per their requirement. With our analysis, we establish that the VLM performances aren’t highly correlated, and models indeed have different strengths and weaknesses. They perform well in some categories, and struggle in others. But the struggling VLM’s performance can be compensated by another VLM which is strong in that category, making appropriate selection a crucial step.
@@
In summary, we aim to address the following research questions (RQs): (1) How to compare VLMs for different types of VQA tasks under practical application requirements? (2) How to appropriately evaluate VLM outputs for VQA that align closely with human judgements? (3) As per current state-of-the-art, which VLM is suited for which application, depending on various external usage constraints?
@@
(i) We release a dataset covering VQA tasks classified into three aspects: task types, application domains, and knowledge type, enabling VLM comparison based on different practical requirements.
(ii) We propose GoEval, which is a multimodal evaluation metric based on GPT-4o [30], and aligns closer to human judgements for visual question-answering.
(iii) We analyze 10 variants of 8 state-of-the-art VLMs of different sizes, families, using our framework to compare their performance under practical settings.
(iv) Using our analysis, we make recommendations on best suited VLMs for a given application requirement under different technical and business constraints.
2 Experimental Dataset Creation
@@ 2.1 Source Datasets
To be able to evaluate VLMs on a wide variety of QA tasks, our experimental dataset is created from five standard datasets - VQAv2 [13], OK-VQA [25], A-OKVQA [31], ChartQA [26] and DocumentVQA [27]. VQAv2 is an extensive VQA benchmark, while OK-VQA and A-OKVQA are used primarily for knowledge-based VQA, where the answer to the question doesn’t lie within the scope of the image. ChartQA and DocumentVQA are taken to evaluate the performance of VLMs on questions based on mathematical graphs and charts, and documents, respectively. From the test split of each dataset , we take task instances, sampled randomly without replacement. We include the image, question, and all reference answers. was used as is the minimum size of test set among the five datasets, thus keeping equal contributions from all datasets. Thus, our final experimental set contains task instances, with equal contributions from each dataset.
2.2 Instance Tags
@@
Using the dataset mentioned above only allows us to differentiate based on task types. But there are more ways in which a practical application can be classified. In order to achieve a multi-dimensional comparison that can be helpful for users, we also classify our experimental dataset in two more aspects - application domains, and knowledge types. This is inspired from a recent work [32], but adapted to suit this task. Application domain is the field a task belongs to, like History, Science, Sports, etc., and the knowledge type is a specific type of knowledge required, like Geographical, Common sense, etc. Both of them are different factors. For example, one can show a crest of a Football team, and ask where is the home stadium of this team located. This is a Sports domain question, but requires a Visual and Geographical Knowledge. Thus, both aspects are necessary.
Our initial set of tags are crafted manually, with an aim to achieve a broad spectrum of application domains, and reasoning types. They are specified in Table 1, and cover a wide range of domains and knowledge types. We tag each of our task instances in the dataset with one or more application domains and knowledge types using the method outlined in Figure 2 and discussed below.
@@ 2.3 Generating Instance Tags
Creating the tags for task type is straightforward - we map ChartQA to ‘Chart Understanding’, DocumentVQA to ‘Document Understanding’, A-OKVQA and OKVQA to ‘Knowledge-based Visual Question-Answering’ and VQAv2 to ‘Visual Question-Answering’.
For application domains and knowledge types, we utilize gpt-3.5-turbo [5, 28]. To correctly generate instance tags, we require key features of the image and question, and also need to eliminate less useful information from the image to avoid confusions in tag generation. To achieve this, following a recent work [12], we generate captions of images from VIT-GPT2222https://huggingface.co/nlpconnect/vit-gpt2-image-captioning, describing key information of the image, and object tags, containing all the objects in the image from Azure Computer Vision API333https://westus.dev.cognitive.microsoft.com/docs/services/computer-vision-v3-2/operations/56f91f2e778daf14a499f21b. Using these two as the descriptor of the image, and the question, we prompt gpt-3.5-turbo to get the application domains and knowledge type. The prompt used for this task is given in Table 7 in the Appendix A (in supplementary material).
After this, we post-process the tags by removing entries which don’t belong to the any of the entries in Table 1. If all tags are removed for a task instance, we add "Other" by default. In the last step, we manually go through each of the classifications and ensure that they are correctly tagged, and replacing any erroneous tag. The final set contains questions, images, candidate answers, task type, application domains and knowledge type of all 5725 candidates.
From the instance tags, we remove the tags for which number of instances are less than 300 from our analyses. Please note, we don’t remove the task instances entirely, as they may contain other tags included in the study, but just not consider those tags in reporting our results in Section 4 due to less number of instances. This gives a final set of 5 task types, 14 application domains, and 9 knowledge types, which are shown in Figure 3. We also report the mean, max and std. of number of tags per instance, along with caption length and number of object tags per instance in Table 2.
@@
We use this dataset for rest of the analysis in this paper, and also release it publicly, so that they can be used by the research community in further research. Find the link to access it in footnote of Page 1.
@@ 3 GoEval: A VQA Evaluation Metric
With the shift from extractive to generative models for QA, the limitations of evaluation metrics based on lexical matching have been exposed, particularly when the answers are correct, but not do not appear in the set of gold answers [17]. A recent work [17] evaluated traditional VQA metrics against GPT-based evaluation for open-domain QA, and found the it to be more aligned with human judgements.
Another alternative is to evaluate via NLG metrics like ROUGE [20], METEOR [3], and later use a threshold to determine correct/wrong. But they also suffer from similar limitations, as evaluated from a previous work [24], which also established the GPT-based evaluation are aligned closer to human judgements. A probable solution is using BERTScore [38], which compares the texts in embedding space, focusing more on semantic similarity.
While the above metrics show good amount of human correlation, these metrics are not equipped to handle multimodal settings of visual question answering. To fill this gap, we propose GoEval - a multimodal metric based on GPT-4o, which can be used to evaluate VQA. Similar to existing works [17, 24], we create a prompt, and ask GPT-4o if the generated answer is correct. However, to incorporate the vision modality in making judgements, we also use the image. We use zero-shot evaluation with prompting, using both GPT-4o and GPT-4o-mini.
Formally, we pick a prompt function from the first row of Table 8 of Appendix A (in the supplementary material), and generate a prompt based on question , reference answers set , and candidate answer . Using this prompt and the image, we ask GPT-4o [30, 29] model to ask if it is correct or not. We also compare our prompt with other existing techniques, like without using the image, performing this in reference-less settings, so that we can compare which technique shows highest correlation with human judgements. We don’t compare with traditional methods, as it is already done in the above discussed works [17, 24], and our method without using the image is similar to what is used in [17].
@@ 3.1 Validating GoEval
To validate GoEval, we first generate outputs on our experimental dataset using Gemini-1.5-Pro [33], a state-of-the-art VLM. Then, we manually evaluate all the answers on the validation set, marking for incorrect answer and for correct answer. These are the human evaluation results.
We compare BERTScore precision, recall, F1-score, and six variants of GoEval using GPT-4o and GPT-4o mini [30], with and without using reference answers, and with and without using images with human evaluations in terms of accuracy and Kendall’s Tau. The exact prompts that we use are outlined in Table 8 of Appendix A (in the supplementary material). When we don’t use the image, we also alter the template text by a little to not ask model to refer the image. Results are detailed in Table 3.
@@
From the results, we can clearly see that using GoEval with reference answers, images, together provides best alignment with human judgement. Without using the image (-R, -I), the performance proves to be very weak, depicting that existing metrics that don’t utilize the image will perform poorly on VQA. Moreover, the differences between models are largely influenced by whether they use all available information. For example, performance of GoEval and GoEval-mini with image but without reference only has a difference of , but as soon as the reference answers are added, increased by 20 units. This shows that all these information are necessary for a successful evaluation. When used, GoEval shows high accuracy and Kendall’s Tau correlation with human evaluation, and GoEval-mini marginally outperforms GoEval.

### v3

@@ Abstract
Visual Question-Answering (VQA) has become key to user experience, particularly after improved generalization capabilities of Vision-Language Models (VLMs). But evaluating VLMs for an application requirement using a standardized framework in practical settings is still challenging. This paper aims to solve that using an end-to-end framework. We present VQA360 – a novel dataset derived from established VQA benchmarks, annotated with task types, application domains, and knowledge types, for a comprehensive evaluation. We also introduce GoEval, a multimodal evaluation metric developed using GPT-4o, achieving a correlation factor of 56.71% with human judgments. Our experiments with state-of-the-art VLMs reveal that no single model excels universally, thus, making a right choice a key design decision. Proprietary models such as Gemini-1.5-Pro and GPT-4o-mini generally outperform others, but open-source models like InternVL-2-8B and CogVLM-2-Llama-3-19B also demonstrate competitive strengths, while providing additional advantages. Our framework can also be extended to other tasks111Code and dataset can be found in the following link: https://github.com/neelabhsinha/vlm-selection-tasks-domains-knowledge-type..
Guiding Vision-Language Model Selection for Visual Question-Answering Across Tasks, Domains, and Knowledge Types
Neelabh Sinha1, Vinija Jain2††thanks: Work does not relate to position at Meta., Aman Chadha3††thanks: Work does not relate to position at Amazon. 1Georgia Institute of Technology 2Meta AI 3Amazon GenAI nsinha68@gatech.edu, hi@vinija.ai, hi@aman.ai
@@ 1 Introduction
Visual Question Answering (VQA) Antol et al. (2015) is the task of answering a question about an image correctly. This field has been faced with constant challenges in terms of the nature of the problem. For example, the question can be about the image directly Goyal et al. (2017a); Zhu et al. (2016); Goyal et al. (2017b), or outside the scope of the image with external knowledge Marino et al. (2019); Schwenk et al. (2022). The images can be a photograph, a mathematical chart Masry et al. (2022); Li and Tajbakhsh (2023), a document screenshot Mathew et al. (2021), or more.
@@
Dedicated methods Zhu et al. (2016); Goyal et al. (2017b); Kafle and Kanan (2017) have long existed to solve different challenges in VQA. But, with the advancement of Vision-Language models (VLMs) Chen et al. (2023); Bai et al. (2023); Team et al. (2024); OpenAI et al. (2024); Liu et al. (2023a) in multimodal research, several applications have started adapting them due to their versatility. This is because after pre-training Lin et al. (2024); Wei et al. (2024) on vast multimodal datasets Chen et al. (2015); Thomee et al. (2016); Changpinyo et al. (2021); Masry et al. (2022); Mathew et al. (2021); Marino et al. (2019), VLMs can effectively generalize across various types of images, and can also incorporate external knowledge beyond the image.
@@
But which VLM to utilize for a given VQA-based requirement? The complexity of this stems from two directions - new VLMs being proposed and the diverse nature of tasks in VQA. New VLMs come up every month now Liu et al. (2023a, 2024); Bai et al. (2023); team (2024); Chen et al. (2023, 2024); Wang et al. (2023); Hong et al. (2024), and they differ in their architecture, training data, training strategy, size, etc., possessing different capabilites. In addition, users often face technical and business constraints in terms of compute, memory, cost of inference, data, and regulatory risks, which can favor specific VLMs over others. Second, tasks may vary from types such as Chart Question-Answering Masry et al. (2022) or Document Understanding Mathew et al. (2021), to application domains such as Science, or Sports, and the type of knowledge required, like Geographical Information, Mathematical Reasoning, and beyond. For an application that can fall into one or more such categories, how do you identify the best suited VLM? How to compare them meaningfully? These are gaps in existing literature. Technical reports of VLMs provide benchmarks and comparison, but they are very theoretical and limited.
To bridge the gap in evaluating VLMs on VQA, we propose an end-to-end framework that provides a standardized paradigm for evaluating vision-language models (VLMs) across three key aspects: task type, application domain, and knowledge type. Existing datasets like VQAv2 Goyal et al. (2017a), OK-VQA Marino et al. (2019), and ChartQA Masry et al. (2022) offer task instances for training and evaluation but lack labels for other practical aspects. Our framework addresses this by developing and sharing a dataset VQA360 derived from the above benchmarks, where tasks are also labeled with their application domains and the knowledge types required for successful completion, as illustrated in Figure 1, allowing for a evaluation in practical settings. Each task can have multiple tags for all these aspects. In addition, traditional NLP evaluation metrics have been shown to be poorly correlated with human judgments for generative models Kamalloo et al. (2023); Liu et al. (2023b), an issue that extends to VLM. To address this, we introduce GoEval, a multimodal evaluation metric leveraging GPT-4o OpenAI et al. (2024), which demonstrates superior alignment with human judgment compared to existing metrics. Together, they complement each other to provide an end-to-end, completely multimodal evaluation framework. Our framework evaluates 10 variations of 8 VLMs, accommodating diverse requirements such as open-source, resource-efficient, or privacy-compliant models.
@@
In summary, our aim is to address the following research questions (RQs): (1) How to compare VLMs for different types of VQA tasks in practical settings? (2) How to evaluate those VLM outputs closely with human judgments? (3) As per current SOTA, which VLM is suited for which application, depending on various external constraints?
@@
(i) We release VQA360 - a dataset of VQA tasks with three labeled aspects: task types, application domains, and knowledge type, enabling comparison based on different practical requirements.
(ii) We propose GoEval, which is a multimodal evaluation metric based on GPT-4o OpenAI et al. (2024), and aligns more closely with human judgments for visual question-answering.
(iii) We analyze 10 variants of 8 state-of-the-art VLMs of different sizes and families, using our framework to compare their performance.
(iv) Using our analysis, we make recommendations on the best-suited VLMs for a given application requirement under different constraints.
2 VQA360: A Practical Evaluation Dataset
@@ 2.1 Source Datasets
To be able to evaluate VLMs in a wide variety of QA tasks, VQA360 is created from five standard datasets - VQAv2 Goyal et al. (2017a), OK-VQA Marino et al. (2019), A-OKVQA Schwenk et al. (2022), ChartQA Masry et al. (2022) and DocumentVQA Mathew et al. (2021). VQAv2 is an extensive VQA benchmark, while OK-VQA and A-OKVQA are used primarily for knowledge-based VQA, where the answer to the question does not lie within the scope of the image. ChartQA and DocumentVQA are taken to evaluate the performance of VLMs on questions based on mathematical graphs and charts, and documents, respectively. From the test split of each dataset , we take task instances, randomly sampled without replacement. was used as it is the minimum size of test set among the five datasets. Thus, our final experimental set contains task instances, with equal contributions from each dataset.
@@
2.2 Instance Tags
Using the dataset mentioned above only allows us to differentiate based on task types. But there are more ways in which a practical application can be classified. To enable that, we also classify VQA360 in two more aspects - application domains, and knowledge types. This is inspired from a recent work Sinha et al. (2024), but adapted to suit this task. The application domain is the field a task belongs to, such as history, science, sports, etc., and the knowledge type is a specific type of knowledge required, such as geographic, common sense, etc.
Our initial set of tags is crafted manually, with an aim to achieve a broad spectrum of application domains and reasoning types. They are specified in Table 1, and cover a wide range of domains and knowledge types. We tag each of our task instances in the dataset with one or more application domains and knowledge types using the method described in Figure 2 and discussed in the following.
@@ 2.3 Generating Instance Tags
For creating the tags for task types, we map ChartQA to ‘Chart Understanding’, DocumentVQA to ‘Document Understanding’, A-OKVQA and OKVQA to ‘Knowledge-based Visual Question-Answering’ and VQAv2 to ‘Visual Question-Answering’.
For application domains and knowledge types, we use gpt-3.5-turbo Brown et al. (2020); OpenAI (2023). To correctly generate instance tags, we require key features of the image and question, and also need to eliminate less useful information from the image to avoid confusions. To achieve this, following a recent work Fu et al. (2023), we generate captions of images from VIT-GPT2222https://huggingface.co/nlpconnect/vit-gpt2-image-captioning, and object tags from the Azure Computer Vision API. Using these two as the descriptors of the image and the question, we prompt gpt-3.5-turbo to get the application domains and knowledge type. The prompt used for this task is given in Table 7 in Appendix A.
After this, we post-process the tags by removing entries that do not belong to any of the entries in Table 1. If all tags are removed for a task instance, we add "Other" by default. Finally, we manually go through each of the labels and ensure that they are correctly tagged and replace any erroneous tag. The final set contains questions, images, candidate answers, task type, application domains, and knowledge type of all 5725 candidates.
From the instance tags, we remove the tags for which number of instances are less than 300 from our analyses. Please note, we do not remove the task instances entirely, as they may contain other tags included in the study, but just not consider those tags in reporting our results in Section 4 due to less number of instances. This gives a final set of 5 task types, 14 application domains, and 9 knowledge types, which are shown in Figure 3. We also report statistis of VQA360 in Table 2.
@@
VQA360 allows extended analysis of diverse VQA tasks and allows looking into performance of VLMs from entirely new perspectives. Further, the creation methodology can also be extended for enriching other datasets and creating benchmarks for evaluating VLMs in different settings.
We use VQA360 for rest of the analysis, and also release it publicly (linked in footnote of Page 1), for the research community to utilize in future research. Although it contains instances of existing benchmarks, it is a one-stop benchmark for an extended evaluation with labels for task types, application domains, and knowledge types. We also provide object tags and captions.
@@ 3 GoEval: A VQA Evaluation Metric
Evaluating QA using lexical matching has significant limitations, particularly when correct answers don’t match with set of gold answers Kamalloo et al. (2023). A recent work Kamalloo et al. (2023) evaluated traditional VQA metrics against GPT-based evaluation for open-domain QA, and found it to be more aligned with human judgments.
Another alternative is to evaluate using NLG metrics such as ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), and then use a threshold to determine correct/wrong. However, they also suffer from similar limitations, as evaluated from previous work Liu et al. (2023b). A probable solution is to use BERTScore Zhang et al. (2019), which compares texts in embedding space, focusing more on semantic similarity.
These metrics may be promising, but they are not equipped to handle multimodal settings. To fill this gap, we propose GoEval - a multimodal metric based on GPT-4o, which can be used to evaluate VQA. Similar to existing works Kamalloo et al. (2023); Liu et al. (2023b), we create a prompt, and ask GPT-4o if the generated answer is correct. However, to incorporate the vision modality in making judgments, we also use the image. We use zero-shot evaluation with prompting, using both GPT-4o and GPT-4o-mini, to compare and contrast performance v/s cost trade-offs.
Formally, we pick a prompt function from the first row of Table 8 of the Appendix A, and generate a prompt based on question , the reference answer set and the candidate response . Using this prompt and the image, we prompt GPT-4o OpenAI et al. (2024); OpenAI (2024) to ask if it is correct. We also compare our prompt against two other ways – without using the image, and without reference answers. This is to understand which technique shows highest correlation with human judgments. We don’t compare with traditional methods, as it is already done in the above discussed works Kamalloo et al. (2023); Liu et al. (2023b), and our method without using the image is similar to what is used in Kamalloo et al. (2023).
@@ 3.1 Validating GoEval
To validate GoEval, we first generate outputs on our experimental dataset using Gemini-1.5-Pro Team et al. (2024), a state-of-the-art VLM. Then, we manually evaluate all the answers on the validation set, marking for incorrect answer and for correct answer. These are the results of human evaluation.
We compare BERTScore precision, recall, F1-score, and six variants of GoEval using GPT-4o and GPT-4o mini OpenAI et al. (2024), with and without reference answers, and with and without using images with human evaluations in terms of accuracy and Kendall Tau. The exact prompts that we use are outlined in Table 8 of Appendix A. When we don’t use the image, we also alter the template text by a little to not ask model to refer the image. The results are detailed in Table 3.
@@
From the results, we can see that GoEval with reference answer and image provides the best alignment with human judgment. Without using the image (-R, -I), the performance proves to be very weak, depicting that existing metrics that do not utilize the image will perform poorly on VQA. We believe this is because image context is crucial for VQA, and in the absence of it, the model isn’t able to reason well whether the provided answer is correct or not. Moreover, the differences between models are largely influenced by the amount of information. For example, performance of GoEval and GoEval-mini with image but without reference only has a difference of , but as soon as the reference answers are added, increases by 20 units. This is because the reference answers serve as an extra guidance in addition of the image to determine the correctness of the candidate answer, making the model perform better.
In summary, GoEval shows high accuracy and Kendall’s Tau correlation with human evaluation, and GoEval-mini marginally outperforms GoEval. This complements VQA360 to provide an end-to-end cohesive framework for VLM evaluation in entirely multimodal settings.